OpenStudy (anonymous):

In 1989, a movie ticket cost $3.00. In 1999, a move ticket costs $5.50. Find the average rate of change of movie tickets price in dollars per year. Answer $0.20 $0.25 $4.00 $0.85

OpenStudy (bahrom7893):

(f(b) - f(a)) / (b-a) = (5.50 - 3.00) / (1999-1989) = 2.50/10 = 0.25

OpenStudy (anonymous):

avg are found by div totals here it shud be (3+5.50)/(1999-1989) since the total no of yrs is 1999-1989=10 8.50/10=.85 $.85

OpenStudy (anonymous):

wats the right opt